OpenStudy (anonymous):

Please help me check my work. I needed the equation for the tangent line to the elipse 2x^2 + y^2 = 3 and that was parallel to 2x + y = 5.

OpenStudy (anonymous):

Wait a sec.

OpenStudy (anonymous):

I first wrote it like this: r: (x,y) = (x0,y0) + lambda v, where v gradient f(x0,y0) = 0. gradient of f(x,y) = (4x,2y) gradient of f(x0,y0) = (4x0,2y0)

OpenStudy (anonymous):

Then I did the same with 2x + y = 5 r2: (x,y) = (x1,y1) + t u gradient of f(x,y) = (2,1) gradient of f(x1,y1) = (2,1) u = (-1,2) v is parallel to (2,1), so for some value of h, v = h (-1,2) v = (2y0,-4x0) 2y0 = - h -4x0 = 2h y0 = -h/2 x0 = -h/2 h = -2 y0 = 1 x0 = 1

OpenStudy (anonymous):

r: (x,y) = (1,1) + lambda (2,-4)

OpenStudy (anonymous):

differentiate the elipse : 4x + 2y(dy/dx) = 0 dy/dx = -2x/y the line given is y = -2x + 5 so we need -2x/y = -2 -> x = y if we put x = y in the equation we get 3x^2 = 3 x = +-1 and so y = +-1 so there are two options x=1 , y =1 slope -2 y = 1 -2(x-1) y=1-2x+2 y = -2x +3 x=-1,y=-1 slope -2 y = -1 -2(x+1) y = -1 -2 -2x y = -2x -3

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Is it right?

OpenStudy (anonymous):

there are two tangent lines possible as i wrote : y = -2x +3 y = -2x -3

OpenStudy (anonymous):

Sorry, your other post didn't show

OpenStudy (anonymous):

oh so you see it now ?

OpenStudy (anonymous):

Yes. But is what I did right too? And if it's wrong, why?

OpenStudy (anonymous):

you got those two lines ?

OpenStudy (anonymous):

I got r: (x,y) = (1,1) + lambda (2,-4)

OpenStudy (anonymous):

im not sure that its the same

OpenStudy (anonymous):

Let me check: (x,y) = (1 + 2lambda, 1 - 4 lambda) x = 1 + 2 lambda y = 1 - 4 lambda y = -2x + b = - 2 - 4 lambda + 3 = 1 - 4 lambda b = 3 y = -2x + 3

OpenStudy (anonymous):

I don't get the other line with -3.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

why you got only x,y = 1,1 i got -1,-1 as well

OpenStudy (anonymous):

-1,-1? I'll see about it.

OpenStudy (anonymous):

y0 = -h/2 x0 = -h/2 h = 2 y0 = -1 x0 = - 1 You're right. But then are there infinite lines? I can choose any h.

OpenStudy (anonymous):

Ha! I get it, any other h will be just a multiple of 1 or -1. You're right. Thank you.

OpenStudy (anonymous):

i dont really fully understand what you did there but if it works for you its fine you see we could do this without using any vector analysis

OpenStudy (anonymous):

You see, any other h, will just be a constant times 1 or - 1. The first curve is an ellipse, and therefore it will have to tangents at two simetrical points parallel to 2x + y = 5.

OpenStudy (anonymous):

ok but i would like you to explain (if you want) your method i understand you took the gradient of the line 2x + y = 5 its (2,1) i dont know why then you chose u(-1,2)

OpenStudy (anonymous):

Because the dot product of the gradient times u = 0.

OpenStudy (anonymous):

but then you find a perpendicular line ?

OpenStudy (anonymous):

The gradient is perpendicular to the vector that gives the direction of the tangent line.

OpenStudy (anonymous):

isnt the gradient tangent to the curve ? :|

OpenStudy (anonymous):

No, it's perpendicular to it. It's only tangent if you have the curve of maximum angle.

OpenStudy (anonymous):

well im confused lol i have to think about it .. thank you

OpenStudy (anonymous):

You're welcome, and thank you for helping me out.